Grade 9Algebra

x + 1/x = 2
then x^37 + 1/x^23 = ?
tell me the answer please !!!!
as soon as possible......
difficult question

SAMBHAV MISHRA
8 Years agoGrade 9

1 Answer

Piyush Kumar Behera
ApprovedApproved Tutor Answer8 Years ago
@Sambhav,In the question it is given that 
 x+\frac{1}{x}=2
\frac{x^{2}+1}{x}=2
x^{2}+1=2x
x^{2}-2x+1=0
(x-1)^{2}=0
Hence
x-1=0  is the only solution
so
x=1
So to find the value of
x^{37}+\frac{1}{x^{23}}
we can substitute x=1
so the answer is
1^{37}+\frac{1}{1^{23}}=1
So the answer is 1.Hope it helps.